Simplify:
3^-2
Please show me step by step to get the answer. thanks

Answer by stanbon(75887)

Simplify:
3^-2
---
The 2 means you should square the base.
The "negative" means you should invert the base.
----
You may perform these instructions in either order:
If you 1st "invert" you get:
3^-2 = (1/3)^2
Then you should square the result to get:
= 1/9
==========================
If you 1st "square" you get:
3^-2 = 9^-1
Then you shoud inver to get:
= 1/9
=============
Cheers,
Stan H.